### v1

# Towards Understanding the Working Mechanism of Text-to-Image Diffusion Model

Abstract
Recently, the strong latent Diffusion Probabilistic Model (DPM) has been applied to high-quality Text-to-Image (T2I) generation (e.g., Stable Diffusion), by injecting the encoded target text prompt into the gradually denoised diffusion image generator. Despite the success of DPM in practice, the mechanism behind it remains to be explored. To fill this blank, we begin by examining the intermediate statuses during the gradual denoising generation process in DPM. The empirical observations indicate, the shape of image is reconstructed after the first few denoising steps, and then the image is filled with details (e.g., texture). The phenomenon is because the low-frequency signal (shape relevant) of the noisy image is not corrupted until the final stage in the forward process (initial stage of generation) of adding noise in DPM. Inspired by the observations, we proceed to explore the influence of each token in the text prompt during the two stages. After a series of experiments of T2I generations conditioned on a set of text prompts. We conclude that in the earlier generation stage, the image is mostly decided by the special token [EOS] in the text prompt, and the information in the text prompt is already conveyed in this stage. After that, the diffusion model completes the details of generated images by information from themselves. Finally, we propose to apply this observation to accelerate the process of T2I generation by properly removing text guidance, which finally accelerates the sampling up to 25%+.
1 Introduction
In real-world application, the Text-to-Image (T2I) generation has long been explored owing to its wide applications [46; 31; 32; 2; 15], whereas the Diffusion Probabilistic Model (DPM) [12; 37; 32] stands out as a promising approach, thanks to its impressive image synthesis capability. Technically, the DPM is a hierarchical denoising model, which gradually purifies noisy data from a standard Gaussian to generate an image. In the existing literature [31; 34; 28; 6], the framework of (latent) Stable Diffusion model [31] is a backbone technique in T2I generation via DPM. In this approach, the text prompt is encoded by a CLIP text encoder [30], and injected into the diffusion image decoder as a condition to generate a target image (latent encoded by VQ-GAN in [31]) that is consistent with the text prompt. Though this framework works well in practice, the working mechanism behind it, especially for the text prompt, remains to be explored. Therefore, in this paper, we systematically explore the working mechanism of stable diffusion.
Our investigation starts from the intermediate status of the denoising generation process. Through an experiment (details are in Section 4), we find that in the early stage of the denoising process, the overall shapes of generated images (latent) are already reconstructed. In contrast, the details (e.g., textures) are then filled at the end of the denoising process. To explain this, we notice that the overall shape (resp. semantic details) is decided by low-frequency (resp. high-frequency) signals [9]. We both empirically show and theoretically explain that in contrast to the high-frequency signals, the low-frequency signals of noisy data are not corrupted until the end stage of the forward noise-adding process. Therefore, its reverse denoising process firstly recovers the low-frequency signal (so that overall shape) in the initial stage, and then recovers the high-frequency part in the latter stage.
Following the phenomenons, we investigate the effect of encoded tokens in the text prompt of T2I generation during the two stages, where each token is encoded by an auto-regressive CLIP text encoder. The text prompt has a length of 76, and is enclosed by special tokens [SOS] and [EOS], 111[EOS] contains the overall information in text prompt due to the auto-regressive CLIP text-encoder at the beginning and end of the text prompt, respectively. Therefore, we categorize the tokens into three classes, i.e., [SOS], semantic tokens, and [EOS]. Notably, the special token [SOS] does not contain information, due to the auto-regressive encoding of the text prompt. Thus, our investigations into the influence of tokens will primarily focus on the semantic tokens and [EOS]. Surprisingly, we find that compared with semantic tokens, the special token [EOS] has a larger impact during generation.
Concretely, under a set of collected text prompts, we select 1000 pairs “[SOS] + Prompt $A$ ($B$) + [EOS]A(B)” from it. Then, replace the special token [EOS]A in the text prompt $A$ with [EOS]B from prompt $B$ to observe the generated data under this condition. Interestingly, we find that the generated images are more likely to be aligned with text prompt $B$ (especially for the shape features) instead of $A$, so that [EOS] has a larger impact compared with semantic tokens. Besides that, we further find that the information in [EOS] is already conveyed during the early shape reconstruction stage of the denoising process. Exploring along the working stage of [EOS], we further verify and explain that the whole text prompts (including semantic ones) primarily work on the early denoising process, when the overall shapes of generated images are constructed. After that, the image details are mainly reconstructed by the images themselves. This phenomenon is explained by “first shape then details”, as the injected text prompt implicitly penalizes the generated images to be consistent with it. Therefore, the penalization quickly becomes weak, when the overall shape of image is reconstructed.
Finally, we apply our observations in one practical cases: Training-free sampling acceleration, as the text prompt works in the first stage of denoising process, we remove the textual prompt-related model propagation ($\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\mathcal{C})$ in (3)) during the details reconstruction stage, which merely change the generated images but save about 25%+ inference cost.
We summarize our contributions as follows.
  1. 1.
    We show, during the denoising process of the stable diffusion model, the overall shape and details of generated images are respectively reconstructed in the early and final stages of it.
  2. 2.
    For the working mechanism of text prompt, we empirically show the special token [EOS] dominates the influence of text prompt in the early (overall shape reconstruction) stage of denoising process, when the information from text prompt is also conveyed. Subsequently, the model works on filling the details of generated images mainly depending on themselves.
  3. 3.
    We apply our observation to accelerate the sampling of denoising process 25%+.
2 Related Work
Diffusion Model.
In this paper, our exploration is based on the Stable Diffusion [32], which now terms to be a standard T2I generation technique based on DPM [12; 36], and has been applied into various computer vision domains e.g., 3D [29; 33; 19] and video generation [26; 4]. In practice, the goal of T2I is generating an image that is consistent with a given text injected into the cross-attention module [42] of the image decoder. Therefore, understanding the working mechanism of stable diffusion potentially improves the existing techniques [1]. Unfortunately, to the best of our knowledge, the problem is limited explored, expected in [45; 35], where they similarly observe the low-frequency signals are firstly recovered in the denoising process. However, further explanations for this phenomenon are neglected in these works.
Influence of Tokens.
Understanding the working mechanism of encoded (by a pretrained language model) text prompt [3; 30; 43; 27] helps us understanding T2I generation [38; 16; 21]. For example, [44] finds that in LLM, the first token primarily decides the weights in the cross-attention map, which similarly appeared in the cross-modality text-image stable diffusion model as we observed. [47] explores the influence of individual tokens in counterfactual memorization. However, in the multi-modality models e.g., [30; 18; 17; 23], whereas the textual information interacted with the image in the cross-attention module as in stable diffusion, the working mechanism of tokens interacts with cross-modality data is limited explored, expected in [1]. They find in a single case that the influence of text prompts may decrease during the denoising process, while they do not proceed to study or apply this phenomenon as in this paper. Recently, [48] finds that the cross-attention map between the text prompt and generated images converges during the denoising process, which is also explained by our observations that the information conveyed during the first few denoising steps. Besides that, unlike ours, their observations are lack of theoretical explanation.
3 Preliminaries
We briefly introduce the (latent) stable diffusion model [31], which transfers a standard Gaussian noise into a target image latent that aligns with pre-given text prompts. Here, the generated data space is a low-dimensional Vector-Quantized (VQ) [7] image latent to reduce the computational cost of generation. One may get the target natural image by decoding the generated image latent. In this paper, the original data (image latent) is denoted by $\boldsymbol{x}_{0}$, and the encoded textual prompt (by CLIP text encoder [30]) is represented by $\mathcal{C}$. The noisy data
$\small\boldsymbol{x}_{t}=\sqrt{\bar{\alpha}_{t}}\boldsymbol{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}_{t},$ (1)
is used as input to diffusion model $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}$ trained by
$\small\min_{\boldsymbol{\theta}}\mathbb{E}\left[\left\|\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\mathcal{C},\emptyset)-\boldsymbol{\epsilon}_{t}\right\|^{2}\right],$ (2)
with $0\leq t\leq T$, $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(0,\boldsymbol{I})$ independent of $\boldsymbol{x}_{0}$, $\bar{\alpha}_{t}\to 0$ (resp. $\bar{\alpha}_{t}\to 1$) for $t\to 0$ (resp. $t\to T$). Here, the noise prediction model $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\mathcal{C},\emptyset)$ is constructed by classifier-free guidance [13] with
$\small\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\mathcal{C},\emptyset)=\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\emptyset)+w\left(\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\mathcal{C})-\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\emptyset)\right),$ (3)
where $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\emptyset)$ is an unconditional generative model, and the $w\geq 0$ is guidance scale. As the model is trained to predict noise $\boldsymbol{\epsilon}_{t}$ in $\boldsymbol{x}_{t}$, and $\boldsymbol{x}_{T}$ approximates a standard Gaussian, we can conduct the reverse denoising process (DDIM [36]) transfers a standard Gaussian to target image $\boldsymbol{x}_{0}$
$\small\boldsymbol{x}_{t-1}=\sqrt{\frac{\bar{\alpha}_{t-1}}{\bar{\alpha}_{t}}}\boldsymbol{x}_{t}+\left(\sqrt{\frac{1-\bar{\alpha}_{t-1}}{\bar{\alpha}_{t-1}}}-\sqrt{\frac{1-\bar{\alpha}_{t}}{\bar{\alpha}_{t}}}\right)\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\mathcal{C},\emptyset).$ (4)
Finally, the diffusion model (usually UNet) takes the text prompt as input to the cross-attention module in each basic block  [^2] The model is stacked by such basic blocks sequentially with residual module, self-attention module, and cross-attention module in each of it. of diffusion model with output ${\rm Attention}(Q,K,V)=\mathrm{Softmax}(QK^{\top}/\sqrt{d})V$ ($d$ is dimension of image feature), where $\phi(\boldsymbol{x}_{t})$ is the feature of image, and
$\small Q=W_{Q}\phi(\boldsymbol{x}_{t});K=W_{K}\mathcal{C};V=W_{V}\mathcal{C}.$ (5)
4 First Overall Shape then Details
In this section, we first explore the image reconstruction process of the stable diffusion model. As noted in [1], the generated image’s overall shape is difficult to be alterted in the final stage of the denoising process. Inspired by this observation, and note that the low-frequency and high-frequency signals of image determine its overall shape and details, respectively [9]. We theoretically and empirically verify that the denoising process recovers the low and high-frequency signals in its initial and final stages, respectively, which explains the phenomenon of first overall shape then details”.
4.1 Two Stages of Denoising Process
[…]
Settings (PromptSet).
As in [14], we use 1600 prompts following the template “a {attribute} {noun}”, with the attribute as an adjective of color or texture. We create 800 text prompts respectively under each of the two categories of attributes. Besides that, we add another extra 1000 complex natural prompts in [14] without a predefined sentence template. These prompts consist of the text prompts set (abbrev PromptSet) we used. The classes of nouns, colors, and textures are respectively 230, 33, and 23 in these prompts. In this paper, we generate images under PromptSet by Stable Diffusion v1.5-Base [31]. Finally, without specification, we use 50 steps DDIM sampling [36].
From [40], though stable diffusion generates encoded VQ image latents [7]. These latents preserve semantic information transformed by text prompt through cross-attention module (5). Notably, in the cross-attention module, the pixel is a weighted sum of token embedding with cross-attention map $\mathrm{Softmax}(QK^{\top}/\sqrt{d})$ as weights. The weights reveal the semantic information of token, as they are the correlations between image query $Q$ and textual key $K$. To check the correlation, we visualize the averaged cross-attention map over all layers of model $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}$ under different time steps $t$, from 50 to 1.
[…]
4.2 Frequency Analysis
Figure 2: Figure LABEL:fig:corrupted_data visualizes the completed noisy data and its high-frequency, and low-frequency parts over different time steps, listed from top to bottom. Figures LABEL:fig:norm and LABEL:fig:ratio measure the low/high-frequency signals of $\boldsymbol{x}_{t}$. In Figure LABEL:fig:norm, “Low_Add_Noisy_Data/eps” means the norm of $\sqrt{\bar{\alpha}_{t}}\boldsymbol{x}_{0}^{\rm low}$ and $\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}_{t}^{\rm low}$, vise versa for “High_…”. On the other hand, Figure LABEL:fig:ratio measures the variation ratio of high/low frequency parts of images during the noising/denoising process. For example, “High_Add_Noise” represents $\|\boldsymbol{x}_{t}^{\rm high}-\boldsymbol{x}_{0}^{\rm high}\|/\|\boldsymbol{x}_{0}^{\rm high}\|$ during noising process.
To further explain the above phenomenon, we refer to the frequency-signal analysis. It has been well explored that the low-frequency signals represent the overall smooth areas or slowly varying components of an image (related to the overall shape). On the other hand, the high-frequency signals correspond to the fine details or rapid changes in intensity (related to attributes like textures) [9]. Thereafter, to explain the “first overall shape then details” in the denoising process, it is natural to refer to the variations in frequency signals of images during the denoising process.
Mathematically, suppose the clean data (image latent) $\boldsymbol{x}_{0}$ has $M\times N$ dimensions for each channel with $\boldsymbol{x}_{t}$ defined in (1). Then the Fourier transformation $F_{\boldsymbol{x}_{t}}(u,v)$ (with $u\in[M],v\in[N]$) of $\boldsymbol{x}_{t}$ is
$F_{\boldsymbol{x}_{t}}(u,v)$ $=\frac{1}{MN}\sum\limits_{k=0}^{M-1}\sum\limits_{l=0}^{N-1}\boldsymbol{x}_{t}^{kl}\exp\left(-2\pi\mathrm{i}\left(\frac{ku}{M}+\frac{lv}{N}\right)\right)$ (6)
$=\sqrt{\bar{\alpha}_{t}}F_{\boldsymbol{x}_{0}}(u,v)+\sqrt{1-\bar{\alpha}_{t}}F_{\boldsymbol{\epsilon}_{t}}(u,v),$
where $\sqrt{-1}=\mathrm{i}$, and $\boldsymbol{x}_{t}^{kl}$ is the $(k,l)$ component of $\boldsymbol{x}_{t}$. As we do not now the distribution of $\boldsymbol{x}_{0}$, we explore the $F_{\boldsymbol{\epsilon}_{t}}(u,v)$ in sequel. The result is in the following proposition proved in Appendix A.
Proposition 1.
For all $u\in[M],v\in[N]$, with probability at least $1-\delta$, we have
$\small\left\|F_{\boldsymbol{\epsilon}_{t}}(u,v)\right\|^{2}\leq\frac{1}{\sqrt{MN}}\left(1+\sqrt{8\log{\frac{2MN}{\delta}}}\right).$ (7)
This proposition indicates that under large image size ($MN$), the strength of frequency signals (no matter low or high) of standard Gaussian are equally close to zero. Thus, the frequency signal of $\boldsymbol{x}_{0}$ in noisy data $\boldsymbol{x}_{t}$ is mainly corrupted by the shrink factor $\bar{\alpha}_{t}$ due to (6), instead of the noise in it.
However, as visualized in Figure LABEL:fig:corrupted_data  [^3] Please note that the $t$ in Figure LABEL:fig:corrupted_data and the other parts of this paper refers to the corresponding $t$-th time step of 50-steps DDIM sampling, which corresponds to $20t$ steps in [12], in contrast to high-frequency part of image, the image’s low-frequency parts  [^4] We distinguish low frequency by threshold 20%, i.e., the lowest 20% parts of spectrum are low-frequency. are more robust than the ones of high-frequency. For example, for $t=20$ in Figure LABEL:fig:corrupted_data, the shape of the clock is still perceptible in the low-frequency part of the image.  [^5] This fact is also observed in [35; 45], while they leverage this phenomenon to refine the architecture of UNet, and the rationale behind the phenomenon is unexplored. If this fact is generalized to image latent, then it explains the two stages of generation as observed in Section 4.1. Because the low-frequency parts are not corrupted until the end of the adding noise process. Then, it will be recovered at the beginning of the reverse denoising process.
To investigate this, in Figures LABEL:fig:norm and LABEL:fig:ratio, we plot the averaged results over time steps of variation of low/high-frequency parts in images generated by PromptSet. In these figures, $\boldsymbol{x}_{t}^{\rm low}$ is the low-frequency part of $\boldsymbol{x}_{t}$ and vice-versa for high-frequency part $\boldsymbol{x}_{t}^{\rm high}$. As can be seen, in Figure LABEL:fig:ratio, the behavior of $\boldsymbol{x}_{t}$ is similar under add/de noise processes, and the reconstruction of low-frequency signals is faster than the high-frequency signals. On the other hand, by comparing “Low_…_Data” ( $\|\sqrt{\bar{\alpha}_{t}}\boldsymbol{x}_{0}^{\rm low}\|$) and “High_…_Data” ($\|\sqrt{\bar{\alpha}_{t}}\boldsymbol{x}_{0}^{\rm high}\|$) in Figure LABEL:fig:norm, we observe the strength of high-frequency signals are significantly lower than the low-frequency signals, which seems to be a property adopted from natural image [9]. However, the relationship oppositely holds for Gaussian noise, which is implied by Proposition 1, as the frequency signals of noise $\boldsymbol{\epsilon}_{t}$ under each spectrum are all close to zero, while the high-frequency parts contain 80% spectrum, so that $\boldsymbol{\epsilon}_{t}^{\rm high}$ is larger than the $\boldsymbol{\epsilon}_{t}^{\rm low}$.
These observations explain the phenomenon “first overall shape then details”. Since the low-frequency parts of the image (decide overall shape) are not totally corrupted until the end of the noising process. Thus, they will be firstly recovered during the reverse denoising process, while the phenomenon does not hold for low-frequency parts of the image, as they are quickly corrupted during the noising process, so they will not be recovered until the end of denoising.
5 The Working Mechanism of Text Prompt
We have verified that the denoising process has two stages i.e., “first overall shape then details”. Next, we explore the working mechanism of text prompts during these stages. Our main observations are two fold, 1): The special tokens [EOS] dominate the influence of text prompt. 2): The text prompt mainly works on the first overall shape reconstruction stage of the denoising process.
5.1 [EOS] Contains More Information
In T2I diffusion, the text prompt is encoded by auto-regressive CLIP text encoder, with semantic tokens (SEM) enclosed with special tokes [SOS] and [EOS]. For such three classes of tokens, as the information in these tokens is conveyed by the cross-attention module, we first compute the averaged weights over pixels in the cross-attention map for each class. The weights are computed by taking the average over PromptSet and presented in Figure 3. As can be seen, the weights of [SOS] are significantly larger than the other classes. However, due to the CLIP text encoder is an auto-regressive model, [SOS] does not contain any semantic information. Therefore, we conclude that the influence of [SOS] is mainly adjusting the whole cross-attention map i.e., weights on the other tokens. A similar phenomenon is observed in single-modality LLM [44]. As the information of text prompt is conveyed by semantic tokens and [EOS], we will focus on them instead of [SOS] in the sequel.
As both SEM and [EOS] contain the semantic information in the text prompt, we first explore which of them has larger impact on T2I generation. To this end, we select 3000 pairs of text prompts from PromptSet (2000 pairs follow the template, the other 1000 pairs have complex prompts), where the two text prompts are represented as “[SOS] + Prompt $A$ ($B$) + [EOS]A(B)”. For each pair, we switch their [EOS] to construct the new text prompt pairs as “[SOS] + Prompt $A$ ($B$) + [EOS]B(A)”.
We examine the generated images under these artificially constructed text prompts (namely Switched-PromptSet (S-PromptSet)). We call $A$ from PromptA as “source” and $B$ from [EOS]B as “target” for “[SOS] + Prompt $A$ + [EOS]B”, and vice versa. For the generated images under these prompts, we measure their alignments with the source and target prompts, respectively. The used metrics are the three standard ones in measuring text-image alignment: CLIPScore [30; 10], BLIP-VQA [18; 14], and MiniGPT4-CoT [49; 14] (details are in Appendix B).
Figure 3: Averaged weights in cross-attention map over pixels of three classes of tokens. For each prompt in PromptSet, the result is obtained by taking average over tokens in each class. The final result is the average over PromptSet. Notably, the weights on [SOS] are all larger than 0.9.
Figure 4: Images under prompts from S-PromptSet with switched [EOS]. The objects are consistent with the ones conveyed by [EOS], while some information in semantic tokens is still conveyed.
The results are in Table 1. Surprisingly, the generated images under the constructed text prompts are more likely to be aligned with the target prompt instead of the source prompt. That says, even with prefixed irrelevant semantic tokens, the information contained in [EOS] dominates the denoising process (especially for overall shape) as in Figure 4. Thus, we conclude that the special tokens [EOS] have a larger impact than semantic tokens in prompt during T2I generation. We have two speculations about this phenomenon. 1): Owing to the auto-regressive encoded text prompt, unlike semantic tokens, [EOS] contains complete textual information, so that it decides the pattern of the generated image. 2): The number of [EOS] is usually larger than semantic tokens, as the prompt is enclosed by [EOS] to length 76. An ablation study in Appendix C verifies this speculation.
In summary, our conclusion in this subsection can be summarized as: In T2I generation, the special token [EOS] decides the overall information (especially shape) of the generated image.
Remark 1.
For the generated images under S-PromptSet, we find some information in semantic tokens is also conveyed, especially for the attribute information in it, e.g., “brown” color in the last image of the first row in Figure 4. We explore this in Appendix D and explain this as: unlike noun information, attributes in semantic tokens may not conflict with the contained information in [EOS] (which quickly decides the overall shape of the generated image), so that has potential to be conveyed.
Figure 5: Desnoising process under text prompt with switched [EOS] in $[a,50]$.
Table 1: The alignment of generated image with its source and target prompts. The prompts are constructed with switched [EOS].
Alignment Prompt Source Target
Text-CLIPScore $\uparrow$ 0.2363 0.2758
[…]
Figure 6: Relative text-image alignments (“current minus worst” over “best minus worst”) with source (or target) prompt under switched [EOS] (substitution in [Start Step, 50], Figure 5). Alignments with source and target prompts are respectively solid and dot lines.
5.2 The Text Prompt Mainly Working on the First Stage
[…]
In Figure 6, the alignment with the target prompt slightly decreases, until the “Start Step” of substitution close to 50. This shows that the information in [EOS] has been conveyed during the first few steps of the denoising, which is the overall shape reconstruction stage according to Section 4.
Following the revealed working stage of [EOS], we explore whether the whole text prompt also works in this stage. If so, the T2I generation will only depends on $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\emptyset)$ in (3) for small $t$. To see this, we vary the $w$ in (3) to control the injected information from the text prompt during the denoising process. Concretely, for $a$ as the starting step of removing text prompt, i.e., during $t\in[0,a)$, we use $w=7.5$, and $w=0$ for $t\in[a,50]$, where $a\in[0,50]$. Then, the text prompt only works for $t\in[0,a)$. We generate target images $\boldsymbol{x}_{0}^{50}$ under PromptSet with standard denoising process ($a=50$), and compare them with the ones $\boldsymbol{x}_{0}^{a}$ generated under varied $a\in[0,50]$ (Figure 7). The image-image alignments are measured by standard metrics CLIPScore and $L_{1}$-distance [9]. To eliminate magnitude, we report relative results, i.e., “current minus worst” over “best minus worst”.
The results are in Figure LABEL:fig:text_prompt_influence. During generation, the text information is absence for $t\in[a,50]$, while Figure LABEL:fig:text_prompt_influence indicates that alignments between $\boldsymbol{x}_{0}^{a}$ and target $\boldsymbol{x}_{0}^{50}$ will quickly be small only for large $a$ (from 30 to 50). This shows that only if removing the textual information under large $t$, its influence to generated image is removed. Therefore, we can conclude: The information of text prompt is conveyed during the early stage of denoising process. Therefore, the overall shape of generated image is mainly decided by the text prompt, while the its details are then reconstructed by itself.
Discussion.
Next, let us explain the phenomenon. Technically, in (3), the $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\mathcal{C},\emptyset)$ was proposed to approximate $\nabla_{\boldsymbol{x}}\log{p_{t}(\boldsymbol{x}_{t}\mid\mathcal{C})}/\sqrt{1-\bar{\alpha}_{t}}$ with decomposition ($p_{t}$ is the density of $\boldsymbol{x}_{t}$)
$\small\nabla_{\boldsymbol{x}}\log{p_{t}(\boldsymbol{x}_{t}\mid\mathcal{C})}=\nabla_{\boldsymbol{x}}\log{p_{t}(\boldsymbol{x}_{t})}+\nabla_{\boldsymbol{x}}\log{p_{t}(\mathcal{C}\mid\boldsymbol{x}_{t})}.$ (8)
Comparing (3) and (8), it holds $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\emptyset)\propto\nabla_{\boldsymbol{x}}\log{p_{t}(\boldsymbol{x}_{t})}$  [^6] This can be verified by the training strategy of it in [31], where $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\emptyset)$ is used to predict noise in noisy data without condition injected. and $w(\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\mathcal{C})-\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\emptyset))\propto\log{p_{t}(\mathcal{C}\mid\boldsymbol{x}_{t})}$. From [37], the denoising process (4) aims to maximize log-likelihood $\log{p_{0}(\boldsymbol{x}_{0}\mid\mathcal{C})}$. Then, moving along the direction $\nabla_{\boldsymbol{x}}\log{p_{t}(\mathcal{C}\mid\boldsymbol{x}_{t})}$ (leads to large $\log{p_{t}(\mathcal{C}\mid\boldsymbol{x}_{t})}$) push $\boldsymbol{x}_{t}$ to be aligned with the text prompt $\mathcal{C}$ during the decreasing of $t$. Adding such a moving direction is standard in conditional generation [25; 36; 8; 24]. As shown in Figure LABEL:fig:noise_norm, during the denoising process, $\boldsymbol{x}_{t}$ will gradually to be consistent with $\mathcal{C}$, so that $\nabla\log{p_{t}(\mathcal{C}\mid\boldsymbol{x}_{t})}$ will decrease with $t$. Thereafter, we observe the impact of text prompt conveyed by this term decreases with $t\to 0$. Notably, owing to the quickly reconstructed overall shape of image in Section 4, the generated $\boldsymbol{x}_{t}$ will quickly be consistent with $\mathcal{C}$, so that explain the quickly decreasing $\nabla\log{p_{t}(\mathcal{C}\mid\boldsymbol{x}_{t})}$.
Figure 7: Desnoising with text prompt injected in $[0,a]$.
Figure 8: Figure LABEL:fig:text_prompt_influence is the relative difference “current minus worst” over “best minus worst” under different start step $a$ of Denoising process Figure 7. The last two figures LABEL:fig:noise_norm are per-dimensional norm of unconditional noise $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\emptyset)$ and noise difference $w\left(\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\mathcal{C})-\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\emptyset)\right)$
6 Application
Acceleration of Sampling.
Since the information contained in text prompt is mainly conveyed by the noise prediction with condition $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\mathcal{C})$, we can consider removing the evaluation of after the first few steps of denoising process. This is because the information in text prompt has been conveyed in this stage, and the computational cost can be significantly reduced without evaluating $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\mathcal{C})$.
Therefore, we substitute the noise prediction $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\mathcal{C},\emptyset)$ as
$\small\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\mathcal{C},\emptyset)=\begin{dcases}\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\emptyset)+w\left(\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\mathcal{C})-\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\emptyset)\right)\qquad&a\leq t;\\ \boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\emptyset)\qquad&0\leq t<a.\end{dcases}$ (9)
By varying $a\to T$ in (9), the inference cost is reduced as an evaluation of $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\mathcal{C})$ is saved.
To evaluate the saved computational cost of using noise prediction (9) during inference and the quality of generated data, we consider applying it on two standard samplers DDIM [36] and DPM-Solver [22] on a benchmark dataset MS-COCO [20] in T2I generation. We consider backbone models Stable-Diffusion (SD) v1.5-Base, SD v2.1-Base [31], and Pixart-Alpha [5]. Concretely, we apply noise prediction (9) with varied $a$ to generate 30K images from 30K text prompts in the test set of MS-COCO, for each sampler and backbone model. We compare the difference (measured by $L_{1}$-distance and Image-Level CLIPScore) between the generated images under $a>0$ and $a=0$ (the standard noise prediction). The results are in Table 2, where we also report the Frechet Inception Distance (FID) score [11] under each $a$ to evaluate the quality of generated images.
Figure 9: Desnoising under $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}$ (9). The text prompt is injected in $[a,T]$, instead of $[0,a]$ in Figure 7.
Figure 10: The generated images with 25 steps DPM-Solver under $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}$ in (9) (Figure 9). The textual information is removed during $t\in[0,a]$. With $a\to 25$, the inference cost is decreased.
The Table 2 indicates that proper $a$ in (9) significantly reduces the computation cost during the inference stage without deteriorate the quality of generated images. For example, SD v1.5 with $a=20$ saves 27+computational cost, but generates images close to the baseline method ($a=0$).
Table 2: The difference between images generated under varied $a$ with the ones of $a=0$. The results are averaged over 30K generated images, and saved latency is evaluated on one NVIDIA V100 GPU.
Start point $a$ 0 (baseline) 10 20 30 40 50 0 (baseline) 5 10 15 20 25
Sampler / Backbone DDIM / SD v1.5 DPM-Solver / SD v1.5
Image-CLIPScore $\uparrow$ 1.000 0.998 0.996 0.971 0.838 0.539 1.000 0.999 0.994 0.956 0.798 0.533
$L_{1}$-distance $\downarrow$ 0.000 0.011 0.022 0.043 0.087 0.195 0.000 0.015 0.027 0.050 0.100 0.188
Saved Latency $\uparrow$ 0.00% 7.84% 18.10% 27.18% 35.24% 48.47% 0.00% 8.36% 17.95% 26.11% 34.86% 47.60%
FID $\downarrow$ 13.772 13.770 13.805 14.012 15.048 19.296 14.297 14.286 14.725 14.985 15.860 19.758
Text-CLIPScore $\uparrow$ 31.040 31.001 30.894 30.493 28.172 16.682 30.992 30.891 30.772 30.205 26.843 16.721
Sampler / Backbone DDIM / SD v2.1 DPM-Solver / SD v2.1
Image-CLIPScore $\uparrow$ 1.000 0.999 0.998 0.986 0.902 0.550 1.000 0.999 0.998 0.988 0.901 0.543
$L_{1}$-distance $\downarrow$ 0.000 0.017 0.041 0.077 0.152 0.386 0.000 0.026 0.046 0.083 0.160 0.369
Saved Latency $\uparrow$ 0.00% 8.68% 18.95% 28.16% 36.19% 47.99% 0.00% 8.75% 18.28% 26.24% 35.48% 47.75%
FID $\downarrow$ 13.014 13.011 13.046 13.247 14.242 18.472 13.507 13.500 13.914 14.159 15.015 18.983
Text-CLIPScore $\uparrow$ 31.413 31.405 31.362 31.113 29.717 16.706 31.339 31.326 31.292 31.045 29.653 16.639
Sampler / Backbone DDIM / Pixart-Alpha DPM-Solver / Pixart-Alpha
Image-CLIPScore $\uparrow$ 1.000 0.999 0.935 0.744 0.643 0.522 1.000 0.999 0.993 0.911 0.648 0.625
$L_{1}$-distance $\downarrow$ 0.000 0.024 0.058 0.103 0.169 0.247 0.000 0.013 0.022 0.046 0.098 0.199
Saved Latency $\uparrow$ 0.00% 8.24% 17.98% 27.20% 34.95% 49.15% 0.00% 7.92% 15.77% 25.18% 33.80% 48.10%
FID $\downarrow$ 22.651 22.884 23.258 25.485 29.760 36.525 18.669 18.520 18.798 19.358 20.494 26.159
Text-CLIPScore $\uparrow$ 28.157 27.979 25.719 19.640 14.986 14.586 30.733 30.721 30.745 29.416 20.629 14.928
7 Conclusion
In this paper, we investigate the working mechanism of T2I diffusion model. By empirical and theoretical (frequency) analysis, we conclude that the denoising process firstly constructs the overall shape then details of the generated image. Next, we explore the working mechanism of text prompts. We find its special token [EOS] has a significant impact on the overall shape in the first stage of the denoising process, in which the information in the text prompt is conveyed. Then, the details of images are mainly reconstructed by themselves in the latter stage of generation. Finally, we apply our conclusion to accelerate the inference of T2I generation, and save 25%+ computational cost.
Appendix A Proofs of Proposition 1
See 1
Proof.
Note that $\boldsymbol{\epsilon}_{t}^{kl}$ (abbreviated as $\boldsymbol{\epsilon}^{kl}$) are i.i.d. Gaussian random variable for each of $k,l$. Thus we have
$F_{\boldsymbol{\epsilon}}(u,v)=\frac{1}{MN}\sum\limits_{k=0}^{M-1}\sum\limits_{l=0}^{N-1}\boldsymbol{\epsilon}^{kl}\exp\left(-2\pi\mathrm{i}\left(\frac{ku}{M}+\frac{lv}{N}\right)\right)=\frac{1}{MN}\sum\limits_{k=0}^{M-1}\sum\limits_{l=0}^{N-1}\boldsymbol{\epsilon}^{kl}\exp\left(-\mathrm{i}\theta^{kl}_{uv}\right),$ (10)
with $\theta^{kl}_{uv}$ is the $kl$-th angle in complex value space, and we may simplify it as $\theta^{kl}$ for ease of notations.
Next, we will show the proposition is a direct consequence of the concentration inequality of Gaussian distribution. We prove our results under one-dimensional Fourier transformation under dimension $M$, where the proof can be easily generalized to a two-dimensional case. Owing to the definition of the norm of complex value, for any specific $u$,
$\left\|F_{\boldsymbol{\epsilon}}(u)\right\|^{2}=F_{\boldsymbol{\epsilon}}(u)\overline{F_{\boldsymbol{\epsilon}}(u)}=\frac{1}{M^{2}}\boldsymbol{\epsilon}^{\top}\Lambda\textbf{1}\textbf{1}^{\top}\bar{\Lambda}\boldsymbol{\epsilon},$ (11)
where $\Lambda=\mathsf{diag}(e^{-i\theta^{0}},\cdots,e^{-i\theta^{M-1}})$. Then let $\boldsymbol{P}=(\sqrt{1/M}\textbf{1}^{\top},\cdots,)^{\top}\bar{\Lambda}$, where $(\sqrt{1/M}\textbf{1}^{\top},\cdots,)^{\top}$ is constructed by vector $\sqrt{1/M}\textbf{1}$ and its orthogonal complement. We can verify that $\boldsymbol{P}$ is an orthogonal matrix. Then, let $\boldsymbol{y}=\boldsymbol{P}\boldsymbol{\epsilon}$, so that $\boldsymbol{y}$ has the same distribution with $\boldsymbol{\epsilon}$. Thus
$\small\frac{1}{M^{2}}\boldsymbol{\epsilon}^{\top}\Lambda\textbf{1}\textbf{1}^{\top}\bar{\Lambda}\boldsymbol{\epsilon}=\frac{1}{M^{2}}\boldsymbol{y}^{\top}\boldsymbol{P}\Lambda\textbf{1}\textbf{1}^{\top}\bar{\Lambda}\bar{\boldsymbol{P}}^{\top}\bar{\boldsymbol{y}}=\frac{1}{M}\boldsymbol{e}_{1}^{\top}\boldsymbol{y}\bar{\boldsymbol{y}}^{\top}\boldsymbol{e}_{1}=\frac{1}{M}(\boldsymbol{y}^{1})^{2},$ (12)
where $\boldsymbol{y}^{1}$ is a standard Gaussian. Thus, by the Berstein’s inequality to sub-exponential random variable i.e., $\chi_{1}^{2}$, we have
$\small\mathbb{P}\left(\left|F_{\boldsymbol{\epsilon}}(u)-\mathbb{E}\left[F_{\boldsymbol{\epsilon}}(u)\right]\right|\geq\delta\right)=\mathbb{P}\left(\left|(\boldsymbol{y}^{1})^{2}-\mathbb{E}\left[\left(\boldsymbol{y}^{1}\right)^{2}\right]\right|\geq\delta\right)\leq 2\exp\left(-\frac{1}{8}\min\left\{\delta^{2},\delta\right\}\right).$ (13)
Since $\delta\in(0,1)$. Thus with probability at least $1-\delta/M$, we have
$\small\frac{1}{M}-\frac{1}{M}\sqrt{8\log{\frac{2M}{\delta}}}\leq\left\|F_{\boldsymbol{\epsilon}}(u)\right\|^{2}\leq\frac{1}{M}+\frac{1}{M}\sqrt{8\log{\frac{2M}{\delta}}},$ (14)
which proves our conclusion. ∎
Appendix B Text-Image Alignment Metrics
In this paper, we mainly use three metrics as in [14] to measure the alignment between text prompt condition and generated image. Next, we give a brief introduction to the these metrics.
CLIPScore (Text).
After extracting the features of generated images and text prompt respectively by CLIP encoder [30], CLIPScore (Text) is the cosine similarity between the two features. Similarly, the CLIPScore (Image) is the cosine-similarity between two image features.
BLIP-VQA.
To improve the limited details capturing capability of the CLIP encoder, [14] propose BLIP-VQA which leverages the visual question answering (VQA) ability of BLIP model [18]. They compare the generated with the target text prompt separately described by several questions. For example, the prompt “A blue bird” can be separated into questions “a bird ?”, “a blue bird ?” etc. Then BLIP-VQA outputs the probability of “Yes” when comparing generated images and these questions.
MiniGPT4-CoT.
The MiniGPT4-CoT [14] combines a strong multi-modality question answering model MiniGPT-4 [49] and Chain-of-Thought [43]. The metric is computed by feeding the generated images to MiniGPT-4, then sequentially asking the model two questions “describe the image” and “predict the image-text alignment score”. By constructing such CoT, the multi-modal model will not ignore the details in generated images.
Appendix C Ablation Study on Number of [EOS]
Figure 11: Relative BLIP-VQA Combine color and texture add complex, CLIP score, MiniGPT-CoT with source (or target) prompt under different number of [EOS]. Here the y-axis is the current BLIP-VQA, CLIP, MiniGPT-CoT over the maximum ones, which is used to alleviate the bias brought by the metric itself.
In this section, we conduct an ablation on the number of [EOS]. Concretely, for each text prompt in S-PromptSet, we repeat the semantic tokens in the text prompt e.g., “[SOS] + a yellow cat a yellow … + [EOS]”, so that the number of [EOS] is reduced in these constructed text prompts. Then, we generate images under these reconstructed text prompts and compare the text-image alignments between generated images and source or target prompts.
The results are in Figure 11. As can be seen, with the increasing of semantic tokens (so that decreasing of [EOS]), the generated images tend to be consistent with the source prompt, instead of the target prompt. Therefore, we speculate that the domination of [EOS] may be partially originated from its larger number, compared with semantic tokens. On the other hand, we observe that [EOS] in the forward positions have larger impacts compared to the latter ones, as the alignments between generated images with target prompts significantly decreased along the x-axis from right to left, in Figure 11. This trends further indicate that the [EOS] may contain more information compared with the latter ones, which indicates the domination of [EOS] originates from their larger number, but also the more information in the first few [EOS].
Appendix D Conveyed Information in Semantic Tokens
During our discussion in Section 5, we conclude that the [EOS] has a larger impact than the ones of semantic tokens during T2I generation. However, as observed in Figure 4, under text prompt with switched [EOS], some information in semantic tokens are still conveyed in the generated images, e.g., blue color in the last image of the first row in Figure 4. Therefore, we explore how this information is conveyed in this section, which also reveals the working mechanism of text prompt.
Firstly, in Figure 12, we visualize the cross-attention map of each tokens under text prompt from S-PromptSet, similar to Figure LABEL:fig:cross-attention_map_ratio. Surprisingly, we find that in cross-attention map of semantic tokens and [EOS] are all visually similar to the shape of the final generated images. The similarity is reasonable for [EOS] as it contains the overall information, so that it is perceptible and transfer their information according to constructed cross-attention map.
On the other hand, when semantic tokens convey their information according to the similar cross-attention, for attributes (color or texture), unlike object information, they are potentially not contradict to overall shape decided by [EOS]. Thus, the information of attributes is more likely to be conveyed in its corresponding pixels. However, this does hold for object/noun tokens whose information is very likely related to shape, which has already been decided by [EOS].
This discussion explains the phenomenon of information in semantic tokens are appeared in the generated images under prompt from S-PromptSet. Combining the observations to the working stage of text prompt in Section 5, we can conclude that the semantic tokens also work in the T2I generation, though it has less impact compared to [EOS].
Figure 12: The visualization of cross-attention map under text prompt with switched [EOS] from S-PromptSet. The pixels corresponding to semantic tokens are in the shape of the final generated data as in text prompt $B$ provide [EOS]. For example, the token “chair” corresponds to pixels in the shape of paint, so its information can not be conveyed, while this phenomenon does exist in the attribute token “leather”.
Figure 13: Generated images with zero or random vectors substitution.
Appendix E More Evidences on [EOS] Contains More Information
Table 3: The alignment of generated results image under different constructed text prompt sets. Here “Sem + EOS” is the original text prompt, and serves as baseline here. Besides that, the CLIPScore (Image) is the image-level alignment of generated images with the ones under “Sem + EOS”.
[…]
Sem + Zero 0.2407 0.6732 0.5392 0.6757
Sem + Rand 0.2153 0.6038 0.3606 0.5428
Zero + EOS 0.2999 0.8887 0.7467 0.6982
Rand + EOS 0.3008 0.8791 0.7669 0.7000
Sem + EOS 0.3110 1.0000 0.8999 0.7412
In this section, we further verify that the impact of [EOS] is larger than the ones of semantic tokens in T2I generation. To further verify this conclusion, under the text prompts with the format of “[SOS] + Sem + [EOS]” from our dataset PromptSet, we substitute all semantic tokens or [EOS] with zero vectors or random Gaussian noise. As a result, we get the 4 sets of text prompts, i.e., “[SOS] + Sem + Zero” (abbrev Sem + Zero), “[SOS] + Sem + Random” (Sem + Rand), “[SOS] + Zero + [EOS]” (Zero + EOS), and “[SOS] + Random + [EOS]” (Rand + EOS). These constructed text prompts ideally contain complete semantic information, and we verify the alignment of the generated images with the corresponding text prompt conditions. The alignments are measured by text-image alignment metrics: CLIPScore [30, 10], BLIP-VQA [18, 14], and MiniGPT4-CoT [49, 14].
The results are summarized in Table 3. As can be seen, as expected for baseline combination “Sem + EOS”, the alignments under text prompts with “EOS” preserved are significantly better than the ones with “Sem” preserved. Thus, the observations further verify our conclusion that the [EOS] has larger influences than semantic tokens during the denoising process.
Moreover, we find the generation is somehow robust, as involving random noise in text prompts still generates semantic meaningful images. We visualize some generated images under constructed text-prompts from Table 3 in Figure 13, which indicates the images under “Zero + EOS” indeed visually have the best quality in alignment, so that consist with Table 3. Besides that, the other combinations generate semantic meaningful images as well, expected for “Sem + Rand” Thus semantic tokens do not contain enough information for generation.
Appendix F Key or Value Dominates the Influence?
Figure 14: Generated examples of Key or Value substitution.
Figure 15: The averaged KL-divergence over pixels and layers.
Table 4: The alignment of generated image with its source and target prompts, under switched [EOS] on Key or Value substitution. Here $KV$-Sub is the complete substitution as in Section 5, which serves as a baseline here.
Alignment Attribute $K$-Sub $V$-Sub $KV$-Sub
Source Target Source Target Source Target
CLIPScore (Text)$\uparrow$ 0.3132 0.1875 0.2538 0.2628 0.2363 0.2758
[…]
Figure 16: More generated examples under tokens from S-PromptSet.
As mentioned in Section 3, the information from [EOS] is conveyed by the cross-attention module. More concretely, the Key ($K$) and Value ($V$) in it, respectively decide the weights and features in the output of the cross-attention module (a weighted sum of features). Next, we explore their individual influence for [EOS] to further reveal the working mechanism of it.
Concretely, as in Section 5, we generate images under the constructed text prompt set S-PromptSet. However, the substitution of [EOS] is only conducted on computing Key or Values in cross-attention module, which are respectively denoted as $K$-Substitution (Sub) and $V$-Sub. For such two substitutions, similar to Table 1, we compare the image-text alignment of generated images with source and target prompts.
The results are summarized in Table 4. As can be seen, substituting the [EOS] in $V$ has a larger influence than the substitutions in $K$. To explain this, as we have observed in Figure 3, the weights on semantic tokens and [EOS] are significantly smaller than the ones on [SOS]. However, the $K$ of [EOS] is only related to these small weights, which have limited influence. In contrast to $K$, the $V$ of [EOS] contains information on features, which can be directly conveyed in generated images. So that we can conclude the value of [EOS] dominates the influence in generation as observed in Table 4. We present some generated images under text prompts from S-PromptSet, but with only Key or Value substituted as in Table 4. The generated images are in Figure 14.
We further verify the variation of the cross-attention map after substituting [EOS]. For each pixel, the cross-attention map of it is a discrete probability distribution. Thus, we compute the KL-divergence [41] between probability distributions under substituted/unsubstituted $K$. The averaged KL-divergence over pixels and layers in models under different denoising steps is presented in Figure 15, where we add the KL-divergence of cross-attention map distribution between a uniform distribution as a baseline. The result shows that even with substituted [EOS], the cross-attention map does not vary much. We speculate that this is because as in Figure 3, the weights in [SOS] dominate the cross-attention map. Thus, in the cross-attention module, altering $K$ has a slighter influence compared with altering $V$.
Appendix G More Generated Images
In this section, we present more generated images.
G.1 [EOS] Substitution
In this subsection, we first present the generated images under text prompts from S-PromptSet in Figure 16. As can be seen, most overall shape of generated images are consistent with the ones conveyed by [EOS].
G.2 Generated Images in Paragraph “Acceleration of Sampling” of Section 6
In this subsection, we present more generated images under noise prediction (9) with varied $a$. The results are in Figure 17 and 18.
Figure 17: The generated images with 50 steps DDIM under $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}$ in (9), where the textual information are $\mathcal{C}$ removed during time steps $t\in[0,a]$. With $a\to 50$, the inference cost is decreased.
Figure 18: More generated images with 25 steps DPM-Solver under $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}$ in (9), where the textual information are $\mathcal{C}$ removed during time steps $t\in[0,a]$. With $a\to 25$, the inference cost is decreased.